# Measuring Progress in Dictionary Learning for Language Model Interpretability with Board Game Models

### v1

@@ 8 Conclusion
Finally, we propose the $p$-annealing method, a modification to the SAE training paradigm that can be combined with other SAE methodologies and results in an improvement in both coverage and board reconstruction.
@@ Acknowledgments
Appendix A Sparse Autoencoder Training Parameters
@@
Appendix B List of Board State Properties
Table 3 summarizes the high-level board state properties considered in $\mathcal{G}_{\text{strategy}}$. The selection of concepts was inspired by McGrath et al. (2022). The column indicated by ##\## denotes the number of individual BSPs per concept. A single BSP per concept indicates we match this condition globally for any corresponding piece.
@@ Appendix B List of Board State Properties
pin 1 Indicates whether there is a pin on the board, such that a player’s piece cannot move without exposing the kind behind it to capture.
@@
Appendix C Performance of Linear Probes and SAEs on Board State Properties
In Figure 3, we present a mean coverage score over strategy board state properties $\mathcal{G}_{\text{BSP}}$. Properties within $\mathcal{G}_{\text{BSP}}$ vary significantly in complexity. For example, queen detection can be inferred directly from the move history, while fork detection requires an accurate representation of the board state. Table 4 shows that linear probe F1-score is below 0.95 for 6 out of 15 properties in $\mathcal{G}_{\text{BSP}}$. This suggests that chess-GPT (Karvonen, 2024) does not represent these properties linearly. Additional experiments are required to determine whether the representation is present at all.
For the board state case, reconstruction is significantly higher than coverage. This is because there are many SAE features that are high precision classifiers for a configuration of squares, such as ”white pawn on e4, white knight of f3”. In cases where coverage is higher than reconstruction (such as for can_check), it is because there are not many features that are over 95% precision for “there is a check move available” from which we can recover if there is an available check move. Coverage is significantly higher because there is at least one feature that has an $F_{1}$-score of 0.54 for can_check, which may not have a precision greater than 95%.
@@ Appendix C Performance of Linear Probes and SAEs on Board State Properties
Appendix D Model Internal Board State Representation
D.1 Othello Models
Previous research of Othello-playing language models found that the model learned a nonlinear model of the board state (Li et al., 2023a). Further investigation found a closely related linear representation of the board when probing for ”my color” vs. ”opponent’s color” rather than white vs. black (Nanda et al., 2023). Based on these findings, when measuring the state of the board in Othello, we represent squares as (mine, yours) rather than (white, black).
D.2 Chess Models
Similar to Othello models, prior studies of chess-playing language models found the same property, where linear probes were only successful on the objective of the (mine, yours) representation and were unsuccessful on the (white, black) representation (Karvonen, 2024). They measured board state at the location of every period in the Portable Game Notation (PGN) string, which indicates that it is white’s turn to move and maintain the (mine, yours) objective. Some characters in the PGN string contain little board state information as measured by linear probes, and there is not a clear ground truth board state part way through a move (e.g., the “f” in “Nf3”). We follow these findings and measure the board state at every period in the PGN string.

### v2

@@ 8 Conclusion
Finally, we propose the $p$-annealing method, a modification to the SAE training paradigm that can be combined with other SAE methodologies and results in an improvement in both coverage and board reconstruction over the Standard SAE architecture.
@@ Acknowledgments
Appendix A Improving and evaluating sparse autoencoders
Despite the success of SAEs at extracting human-interpretable features, they fail to perfectly reconstruct the activations Cunningham et al. [2023]. One challenge in the training of SAEs with an $L_{1}$ penalty is shrinkage (or ’feature suppression’); in addition to encouraging sparsity, an $L_{1}$ penalty encourages feature activations to be smaller than they would be otherwise. Wright and Sharkey [2024] approached this problem by fine-tuning the sparse autoencoder without a sparsity penalty. Appendix E further quantifies shrinkage across a suite of SAEs trained on chess and Othello models. Jermyn et al. [2024] and Riggs and Brinkmann [2024] explored alternative sparsity penalties to reduce feature suppression during training. Rajamanoharan et al. [2024] introduced Gated SAEs, an architectural variation for the encoder which both addresses shrinkage and improves on the Pareto frontier of $L_{0}$ vs reconstruction error. Recently, Gao et al. [2024] systematically evaluated the scaling laws with respect to sparsity, autoencoder size, and language model size.
The goal of dictionary learning in machine learning is to produce human-interpretable features and capture the underlying model’s computations Bricken et al. [2023]. However, quantitatively measuring interpretability is difficult and often involves manual inspection. Therefore, most existing work assesses the quality of SAEs along different proxy metrics: (1) The cross-entropy loss recovered, which reflects the degree to which the original loss of the language model can be recovered when replacing activations with the autoencoder predictions. (2) The $L_{0}$-norm of feature activations $\mathbb{E}_{z\sim\mathcal{D}}\left\|h(z)\right\|_{0}$, measuring the number of activate features given an input Ferrando et al. [2024]. Makelov et al. [2024] proposed to compare SAEs against supervised feature dictionaries in a natural language setting. However, this requires a significant understanding of the model’s internal computations and is thus not scalable.
Appendix B Sparse Autoencoder Training Parameters
@@
Appendix C List of Board State Properties
Table 3 summarizes the high-level board state properties considered in $\mathcal{G}_{\text{strategy}}$. The selection of concepts was inspired by McGrath et al. [2022]. The column indicated by ##\## denotes the number of individual BSPs per concept. A single BSP per concept indicates we match this condition globally for any corresponding piece.
@@ Appendix C List of Board State Properties
pin 1 Indicates whether there is a pin on the board, such that a player’s piece cannot move without exposing the king behind it to capture.
@@
Appendix D Performance of Linear Probes and SAEs on Board State Properties
In Figure 3, we present a mean coverage score over strategy board state properties $\mathcal{G}_{\text{BSP}}$. Properties within $\mathcal{G}_{\text{BSP}}$ vary significantly in complexity. For example, queen detection can be inferred directly from the move history, while fork detection requires an accurate representation of the board state. Table 4 shows that linear probe F1-score is below 0.95 for 6 out of 15 properties in $\mathcal{G}_{\text{BSP}}$. This suggests that chess-GPT Karvonen [2024] does not represent these properties linearly. Additional experiments are required to determine whether the representation is present at all.
For the board state case, reconstruction is significantly higher than coverage. This is because there are many SAE features that are high precision classifiers for a configuration of squares, such as "white pawn on e4, white knight of f3". In cases where coverage is higher than reconstruction (such as for can_check), it is because there are not many features that are over 95% precision for “there is a check move available” from which we can recover if there is an available check move. Coverage is significantly higher because there is at least one feature that has an $F_{1}$-score of 0.54 for can_check, which may not have a precision greater than 95%.
@@ Appendix D Performance of Linear Probes and SAEs on Board State Properties
Appendix E Relative Reconstruction Bias
Training Standard SAEs with an $L_{1}$ penalty, as described in Section 4, causes a systematic underestimation of feature activations. Wright and Sharkey Wright and Sharkey [2024] term this phenomenon shrinkage. Following Rajamanoharan et al. Rajamanoharan et al. [2024], we measure the relative reconstruction bias $\gamma$ of our SAEs, defined as:
$\gamma:=\arg\min_{\gamma^{\prime}}\mathbb{E}_{x\sim\mathcal{D}}\left[\|\hat{x}_{\text{SAE}}(x)/\gamma^{\prime}-x\|_{2}^{2}\right]$ (15)
Here, $\mathcal{D}$ denotes a large dataset of model internal activations. Intuitively, $\gamma<1$ indicates shrinkage. A perfectly unbiased SAE would have $\gamma=1$.
Our experiments show that p-annealing achieves similar relative reconstruction bias improvements to gated SAEs, both outperforming the baseline architecture. Figure 5 shows that improvements manifest differently across domains: Chess SAEs show a narrower range of bias ($\gamma\approx 0.98$) compared to Othello ($\gamma\approx 0.80$). This domain-dependent variation may reflect differences in the underlying models or data distributions. We observe unstable $\gamma$ values for SAEs with L0 near zero, which represent degenerate cases outside the typical operating range of these models.
(a) Coverage of Board State
(b) Coverage of Board State vs. $L_{0}$
Figure 5: Comparison of the relative reconstruction bias metric $\gamma$ quantifying feature activation shrinkage across a suite of SAEs. $\gamma<1$ indicates shrinkage. A perfectly unbiased SAE would have $\gamma=1$.
Appendix F Model Internal Board State Representation
F.1 Othello Models
Previous research of Othello-playing language models found that the model learned a nonlinear model of the board state [Li et al., 2023a]. Further investigation found a closely related linear representation of the board when probing for "my color" vs. "opponent’s color" rather than white vs. black [Nanda et al., 2023]. Based on these findings, when measuring the state of the board in Othello, we represent squares as (mine, yours) rather than (white, black).
F.2 Chess Models
Similar to Othello models, prior studies of chess-playing language models found the same property, where linear probes were only successful on the objective of the (mine, yours) representation and were unsuccessful on the (white, black) representation [Karvonen, 2024]. They measured board state at the location of every period in the Portable Game Notation (PGN) string, which indicates that it is white’s turn to move and maintain the (mine, yours) objective. Some characters in the PGN string contain little board state information as measured by linear probes, and there is not a clear ground truth board state part way through a move (e.g., the “f” in “Nf3”). We follow these findings and measure the board state at every period in the PGN string.
@@ F.2 Chess Models
Appendix G Additional examples of learned SAE features
We present two additional examples of learned SAE features that we (subjectively) match to board state properties based on their maximally activating input PGN-strings in Figure 6.
(a) "En passant available" detector learned by an SAE.
(b) "Knight on e2 or e7" detector learned by an SAE.
(c) Another example of the "Knight on e2 or e7" detector shown in subfigure (b) above. We interpret this feature as representing a mirrored perspective. It may be firing for "opponent knight on e column, 1 square away from opponent back rank", which is why it fires for both e2 (if black’s turn to move) and e7 (if white’s turn to move).
Figure 6: Additional examples of learned SAE features. We show the full board state of a chosen game in which the SAE latent has a high activation. The PGN-string (model input) which represents the game history is shown below the board. Tokens that activate SAE features are marked in blue, where darker shades correspond to higher feature activations. Moves that create the considered a board state are highlighted in yellow.